OpenStudy (anonymous):

Expand. Express in terms of sums and differences loga^5x^5yz^6

OpenStudy (anonymous):

I'm assuming everything is inside the log function?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I'd start expanding with \[\ln (ab)=\ln a+\ln b\]

OpenStudy (anonymous):

\[loga ^{5x5}yz ^{6}\]

OpenStudy (anonymous):

And then I'd use \[\ln x^a = a \ln x\]

OpenStudy (anonymous):

\[5x5\] is in the exponent?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the log of a product is the sum of logs of each term in the product, so you can split that up, like i said with \[\log (ab) = \log a+\log b\]

OpenStudy (anonymous):

and you can bring the exponent in a log function down to be a product with the log of the base, like i said with \[\log a^n = n\times \log a\]

OpenStudy (anonymous):

Unless I've missed something, you can fully expand the given function using only those two laws of logarithms